Crossett Trucking Company claims that the mean weight of its delivery trucks when they are fully loaded is 6,050 pounds and the standard deviation is 320 pounds. Assume that the population follows the normal distribution. Fifty-five trucks are randomly selected and weighed. Within what limits will 99 percent of the sample means occur?
Answer #1

Given mean weight CH) = 6050 Standard deviation Co-)- 320 Randomly selected difty-five trucke n = 55 Answer: p(x, <B>X) = 0.9x 11 -ke orn - - 2.58 X-6050 = -2.58 390455 X, - 6050 = -2.58 * 43.1487912 X - - 11-32388) + 6050 = 5938.67612 * *, - 5938.7» X-H - ohn - 2.58 X-6050 3901155 .58 X2 - 6050 = 2.58 * 43.1487912 X2 = 111-323881 + 6050 - 6161.323 88 # X2 = (6161.3 (poun